# Coupled Distributional Random Expert Distillation for World Model Online Imitation Learning

### v1

@@ 1 Introduction
Imitation Learning (IL) has recently shown remarkable effectiveness across a wide range of domains, particularly in addressing complex real-world challenges. In robotics, IL has significantly advanced the state of the art in manipulation tasks [55, 44, 42, 6], enabling robots to perform intricate operations with precision and adaptability. Similarly, IL has achieved impressive results in locomotion tasks [7, 41, 29], where it has facilitated the development of robust and agile motion controllers for various robotic platforms. Beyond robotics, IL has also demonstrated its versatility in domains such as autonomous driving [36, 3, 5], where it is used to model complex decision-making processes and ensure safe and efficient vehicle navigation. Moreover, IL has started making meaningful contributions to healthcare [10], providing support in medical decision-making and enhancing the interpretability of complex diagnostic processes. These achievements highlight the broad applicability of IL and its potential to drive transformative progress across diverse fields.
@@
The simplest approach to imitation learning is to apply behavioral cloning directly to the provided expert dataset, as demonstrated in prior works like IBC [14] and Diffusion Policy [6]. However, this approach is not dynamics aware and may result in lack of generalization when encountering out-of-distribution states. To address these shortcomings, methods like GAIL [27], SQIL [38], IQ-Learn [16], MAIL [2] and CFIL [15] have introduced value or reward estimation to facilitate a deeper understanding of the environment, while leveraging online interactions to enhance exploration. Specifically, GAIL, MAIL, and IQ-Learn frame the imitation learning problem as an adversarial training process, distinguishing between the state-action distributions of the expert and the learner.
@@
Recent advancements in latent world models for imitation learning have made significant progress. Several prior works, including V-MAIL [37], CMIL [32], Ditto [9], EfficientImitate [51], and IQ-MPC [34], have integrated adversarial imitation learning frameworks with world models to address imitation learning tasks. However, as discussed in Section D.7, we found that even with world models, the adversarial objectives can still suffer from instability in certain scenarios. To overcome this issue, we propose replacing the adversarial reward or value formulation with a novel density estimation approach based on random network distillation (RND) [4], which mitigates the instability. Specifically, we perform density estimation in the latent space of the world model, leveraging the superior properties of latent representations and their enhanced dynamics-awareness, as the latent dynamics model is trained directly within this space. Unlike existing methods that use RND for imitation learning [45], our approach jointly learns the reward model and other components of the world model, estimating both the expert and behavioral distributions simultaneously. In contrast, the existing Random Expert Distillation [45] estimates distributions in the original observation and action spaces and decouples the reward model learning from the downstream RL process, making it hard to solve complex tasks with high dimensional observation and action spaces. We evaluate our approach across a range of tasks in DMControl [43], Meta-World [52], and ManiSkill2 [18], demonstrating stable performance and achieving expert-level results.
@@ 2.1 Random Network Distillation
Random Network Distillation (RND) [4] is a technique for promoting exploration. In details, it leverages a fixed randomly parameterized network $f_{\bar{\theta}}(x)$ and a learnable predictor network $f_{\theta}(x)$. During training, RND minimizes the following MSE loss for dataset $\mathcal{D}$ for certain data distribution $\rho$:
@@
During the evaluation, we obtain a data point $x^{\prime}$ for unknown data distribution $\rho^{\prime}$. By computing the L2 norm $\|f_{\bar{\theta}}(x^{\prime})-f_{\theta}(x^{\prime})\|_{2}^{2}$, we can estimate the difference between distribution $\rho$ and $\rho^{\prime}$. This can also be interpreted as performing density estimation for the new data point $x^{\prime}$ within the original distribution $\rho$. A similar methodology has been used in imitation learning and inverse reinforcement learning [45].
@@ 2.2 World Models
Recent world models in the context of robotics control and reinforcement learning often represent a model-based RL method with latent spaces. The model learns a latent state transition model $\mathbf{z}^{\prime}=d_{\theta}(\mathbf{z},\mathbf{a})$, along with a encoder $\mathbf{z}=h_{\theta}(\mathbf{z})$ and a policy model $\mathbf{a}=\pi_{\theta}(\mathbf{z})$. The decision-making process often includes planning with latent unrolling. For models based on the Recurrent State-space Model (RSSM) [22], the latent states often are split into a deterministic part and a stochastic part. PlaNet [21] and Dreamer series [22, 23, 24] leverage decoders for observation reconstruction, while TD-MPC series [25, 26] leverages a decoder-free architecture and conducts planning solely in the latent space.
@@ 3 Methodology
In this section, we will go over the motivation and detailed methodology of our method, Coupled Distributional Random Expert Distillation, or CDRED as an abbreviation. We show that our method is stabler and more reasonable compared to naively apply Random Expert Distillation (RED) [45] on imitation learning with world models.
@@ 3.1 Motivation
Random Expert Distillation (RED) [45] performs imitation learning by estimating the support of expert policy distribution. During training, it minimizes $K$ pairs of predictors and fixed random targets in expert dataset with $N$ data points $\mathcal{D}_{E}=\{\mathbf{s}_{i},\mathbf{a}_{i}\}_{0:N}$:
[…]
@@
To address these difficulties, we introduce a coupled approach. This approach jointly estimates both the expert distribution and the behavioral distribution; it encourages policy exploration during the early stages of training. We provide an intuitive illustration in Figure 2 and describe the detailed methodology in Section 3.3. In this coupled approach, we need to estimate the behavioral distribution during online training, which naturally raises the problem of inconsistent final rewards, as noted by [49]. Thus, we adopt their method for tracking the frequency of data occurrence, which we describe in Section 3.2.
@@ 3.2 Mitigating Inconsistent Reward Estimation
Inconsistencies can arise at various stages of RND model training [49]. During the initial stage, these inconsistencies stem from extreme values in one network, which can be mitigated by using multiple target networks (denoted as $K$ target networks). In the final stage, inconsistencies occur when the resulting reward distribution does not align with the actual state-action distribution. To address this, an unbiased estimator for the state-action occurrence count $n$ is necessary. We should track state-action occurrence frequencies in order to maintain consistency when the distributional RND model is trained online. In this section, we replace the original state $\mathbf{s}_{t}$ with the latent representation $\mathbf{z}_{t}$ for the input of the RND model. Following [49], we denote the random variable $c(\mathbf{z}_{t},\mathbf{a}_{t})$ as the output of a target network $f_{\bar{\theta}_{k}}$, where $k$ is sampled uniformly from the interval $[0,K)$. For a predictor $f$ estimating a distribution $\rho$ (which can be either the expert distribution $\rho_{E}$ or the behavioral distribution $\rho_{\pi}$), by minimizing the $L_{2}$-norm loss $\|f(\mathbf{z}_{t},\mathbf{a}_{t})-c(\mathbf{z}_{t},\mathbf{a}_{t})\|_{2}^{2}$, the optimal predictor $f^{*}(\mathbf{z}_{t},\mathbf{a}_{t})$ is given by:
@@
where $c_{i}(\mathbf{z}_{t},\mathbf{a}_{t})$ is representing the $c(\mathbf{z}_{t},\mathbf{a}_{t})$ for the $i$-th occurrence for state-action pair $(\mathbf{z}_{t},\mathbf{a}_{t})$ in distribution $\rho$. In order to track the occurrence count $n$, we adopt a lemma proposed by [49]:
@@
See Appendix E or prior work [49]. ∎
In this way, we are able to estimate the data distribution with higher consistency as the training proceeds. Following [49], we construct the following estimator for $\sqrt{1/n}$ as an additional bonus correction term:
@@ 3.3 Coupled Distributional Random Expert Distillation
We construct a reward model with two predictor networks that share the same random target ensemble on the latent space of a world model. The distributional random target ensemble consists of $K$ random networks $\{f_{\bar{\theta}_{k}}\}_{0:K}$ with fixed parameters. Regarding the predictors, one of them is the expert predictor $f_{\phi}$ while the other is the behavioral predictor $f_{\psi}$. A predictor $f$ is defined by $f:\mathcal{Z}\times\mathcal{A}\rightarrow\mathbb{R}^{p}$, while $p$ is the dimension of the low-dimensional embedding space for L2 norm distance computation. Following [49], we ask these two predictors to learn the random targets sampled. This is different to RED which learn $K$ predictors for $K$ targets. Given an expert buffer $\mathcal{B}_{E}$ and a behavioral buffer $\mathcal{B}_{\pi}$, we aim to optimize through the following objective:
@@
The scalar coefficient $\alpha$ in Eq. 8 balances the contributions of the first term (the $L_{2}$-norm) and the second term (an estimator for $\sqrt{1/n}$). Following [49], we let the first term dominate initially, switching to the second term as training progresses. This can be achieved with a fixed $\alpha$, rather than a dynamic coefficient. This modification enables consistent online estimation of the state-action distribution, directly supporting reward modeling for online imitation learning.
@@ 3.4 Integrating into World Models for Imitation Learning
World models learn the policy and underlying environment dynamics by encoding the observations into a latent space and learning the transition model in the latent space. Decoder-free world models such as TD-MPC series [25, 26] has proved to be a powerful tool for complex reinforcement learning tasks. We leverage a decoder-free world model containing the following components:
@@
The first term contains consistency loss and temporal difference loss to ensure the prediction consistency of the dynamics model and the accuracy for value function estimation. the temporal difference target is computed by $q_{t}=R(\mathbf{z}_{t},\mathbf{a}_{t})+\gamma Q(\mathbf{z}^{\prime}_{t},\pi(\mathbf{z}^{\prime}_{t}))$ where $R(\mathbf{z}_{t},\mathbf{a}_{t})$ is the output of the reward model. We convert the regression TD objective into a classification problem for stabler value estimation, which is also used by the TD-MPC series and mentioned by [12]. $\text{CE}(\hat{q}_{t},q_{t})$ is the cross entropy loss between target Q value and current predicted value. The second term is the reward loss, which is shown in Eq.6. Similar to the computation of reward loss, we also sum up the consistency and TD loss with factor $\lambda$ over a horizon $H$.
@@
Regarding the policy prior update, we adopt maximum entropy objective [20] to train a stochastic policy:
@@
Following TD-MPC series [25, 26], we also leverage model predictive path integral (MPPI) [48] for planning. We optimize using the sampled action sequences $(\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H})$ in a derivative-free style, maximizing the estimated return for the latent trajectories that have been rolled out using our dynamics model. Mathematically, our objective can be describe as a return maximizing process [26]:
@@ 4 Experiments
We conduct experiments across a diverse range of tasks, including locomotion, manipulation, and tasks with both visual and state-based observations. We evaluate our approach using the DMControl [43], Meta-World [52] and ManiSkill2 [18] environments. As for the baselines, we compare our approach with IQ-MPC [34], which integrates a world model architecture, as well as with model-free approaches, specifically IQ-Learn+SAC [16] (referred to as IQL+SAC in the plots), CFIL+SAC [15], HyPE [39] (In Appendix D.3) and SAIL [46] (In Appendix D.4). Additionally, we also incorporate behavioral cloning (BC) as a baseline method in our evaluation. For all experiments, we sample expert trajectories from a trained TD-MPC2 [26]. Additionally, we conduct ablation studies on the number of expert trajectories and the choice of function $g$, as detailed in Appendix D.2. We further evaluate the robustness of our algorithm in noisy environment dynamics (Appendix D.5), examine the benefits of constructing the reward model in the latent space (Appendix D.8), and highlight its advantages over existing adversarial training methods (Appendix D.7). We also provide the quantitative results measuring the training stability in Appendix D.6. All of the experiments are conducted on a single RTX3090 graphic card.
@@ 4.1 Meta-World Experiments
We conduct experiments on 6 tasks in Meta-World environments. We use 100 expert trajectories for each task, ensuring that the expert data remains consistent across all algorithms for fair comparison within each task. IQ-MPC suffers from overly powerful discriminators in these tasks, even with gradient penalty applied, due to the adversarial training methodology. CFIL+SAC [15] encounters instability in the training process due to the challenges inherent in training flow models. We show stable and expert-level performance, outperforming these baselines in these tasks. We show the episode reward results in Figure 4 and success rate results in Table 1.
@@
Box Close 0.58$\pm$0.12 0.61$\pm$0.09 0.00$\pm$0.00 0.53$\pm$0.18 0.96$\pm$0.03
Bin Picking 0.43$\pm$0.18 0.75 $\pm$ 0.11 0.01$\pm$0.01 0.79$\pm$0.05 0.99$\pm$0.01
Reach Wall 0.10$\pm$0.08 0.90$\pm$0.04 0.01$\pm$0.01 0.31$\pm$0.14 0.98$\pm$0.01
Stick Pull 0.02$\pm$0.02 0.34$\pm$0.11 0.00$\pm$0.00 0.13$\pm$0.08 0.92$\pm$0.05
Stick Push 0.42$\pm$0.14 0.76$\pm$0.14 0.00$\pm$0.00 0.23$\pm$0.10 0.94$\pm$0.03
Soccer 0.04$\pm$0.03 0.73$\pm$0.09 0.01$\pm$0.01 0.12$\pm$0.07 0.81$\pm$0.05

### v2

@@ 1 Introduction
Imitation Learning (IL) has recently shown remarkable effectiveness across a wide range of domains, particularly in addressing complex real-world challenges. In robotics, IL has significantly advanced the state of the art in manipulation tasks [zhu2022viola, wan2024lotus, stepputtis2020language, chi2023diffusionpolicy], enabling robots to perform intricate operations with precision and adaptability. Similarly, IL has achieved impressive results in locomotion tasks [chiu2024learning, seo2023deep, huang2024diffuseloco], where it has facilitated the development of robust and agile motion controllers for various robotic platforms. Beyond robotics, IL has also demonstrated its versatility in domains such as autonomous driving [pan2017agile, bronstein2022hierarchical, cheng2024pluto], where it is used to model complex decision-making processes and ensure safe and efficient vehicle navigation. Moreover, IL has started making meaningful contributions to healthcare [deuschel2023contextualized], providing support in medical decision-making and enhancing the interpretability of complex diagnostic processes. These achievements highlight the broad applicability of IL and its potential to drive transformative progress across diverse fields.
@@
The simplest approach to imitation learning is to apply behavioral cloning directly to the provided expert dataset, as demonstrated in prior works like IBC [florence2022implicit] and Diffusion Policy [chi2023diffusionpolicy]. However, this approach is not dynamics aware and may result in lack of generalization when encountering out-of-distribution states. To address these shortcomings, methods like GAIL [ho2016generative], SQIL [reddy2019sqil], IQ-Learn [garg2021iq], MAIL [baram2016model] and CFIL [freund2023coupled] have introduced value or reward estimation to facilitate a deeper understanding of the environment, while leveraging online interactions to enhance exploration. Specifically, GAIL, MAIL, and IQ-Learn frame the imitation learning problem as an adversarial training process, distinguishing between the state-action distributions of the expert and the learner.
@@
Recent advancements in latent world models for imitation learning have made significant progress. Several prior works, including V-MAIL [rafailov2021visual], CMIL [kolev2024efficient], Ditto [demoss2023ditto], EfficientImitate [yin2022planning], and IQ-MPC [li2024rewardfreeworldmodelsonline], have integrated adversarial imitation learning frameworks with world models to address imitation learning tasks. However, as discussed in Section D.7, we found that even with world models, the adversarial objectives can still suffer from instability in certain scenarios. To overcome this issue, we propose replacing the adversarial reward or value formulation with a novel density estimation approach based on random network distillation (RND) [burda2018exploration], which mitigates the instability. Specifically, we perform density estimation in the latent space of the world model, leveraging the superior properties of latent representations and their enhanced dynamics-awareness, as the latent dynamics model is trained directly within this space. Unlike existing methods that use RND for imitation learning [wang2019random], our approach jointly learns the reward model and other components of the world model, estimating both the expert and behavioral distributions simultaneously. In contrast, the existing Random Expert Distillation [wang2019random] estimates distributions in the original observation and action spaces and decouples the reward model learning from the downstream RL process, making it hard to solve complex tasks with high dimensional observation and action spaces. We evaluate our approach across a range of tasks in DMControl [tassa2018deepmind], Meta-World [yu2020meta], and ManiSkill2 [gu2023maniskill2], demonstrating stable performance and achieving expert-level results.
@@ 2.1 Random Network Distillation
Random Network Distillation (RND) [burda2018exploration] is a technique for promoting exploration. In details, it leverages a fixed randomly parameterized network $f_{\bar{\theta}}(x)$ and a learnable predictor network $f_{\theta}(x)$. During training, RND minimizes the following MSE loss for dataset $\mathcal{D}$ for certain data distribution $\rho$:
@@
During the evaluation, we obtain a data point $x^{\prime}$ for unknown data distribution $\rho^{\prime}$. By computing the L2 norm $\|f_{\bar{\theta}}(x^{\prime})-f_{\theta}(x^{\prime})\|_{2}^{2}$, we can estimate the difference between distribution $\rho$ and $\rho^{\prime}$. This can also be interpreted as performing density estimation for the new data point $x^{\prime}$ within the original distribution $\rho$. A similar methodology has been used in imitation learning and inverse reinforcement learning [wang2019random].
@@ 2.2 World Models
Recent world models in the context of robotics control and reinforcement learning often represent a model-based RL method with latent spaces. The model learns a latent state transition model $\mathbf{z}^{\prime}=d_{\theta}(\mathbf{z},\mathbf{a})$, along with a encoder $\mathbf{z}=h_{\theta}(\mathbf{z})$ and a policy model $\mathbf{a}=\pi_{\theta}(\mathbf{z})$. The decision-making process often includes planning with latent unrolling. For models based on the Recurrent State-space Model (RSSM) [hafner2019learning], the latent states often are split into a deterministic part and a stochastic part. PlaNet [hafner2019dream] and Dreamer series [hafner2019learning, hafner2020mastering, hafner2023mastering] leverage decoders for observation reconstruction, while TD-MPC series [hansen2022temporal, hansen2023td] leverages a decoder-free architecture and conducts planning solely in the latent space.
@@ 3 Methodology
In this section, we will go over the motivation and detailed methodology of our method, Coupled Distributional Random Expert Distillation, or CDRED as an abbreviation. We show that our method is stabler and more reasonable compared to naively apply Random Expert Distillation (RED) [wang2019random] on imitation learning with world models.
@@ 3.1 Motivation
Random Expert Distillation (RED) [wang2019random] performs imitation learning by estimating the support of expert policy distribution. During training, it minimizes $K$ pairs of predictors and fixed random targets in expert dataset with $N$ data points $\mathcal{D}_{E}=\{\mathbf{s}_{i},\mathbf{a}_{i}\}_{0:N}$:
$\hat{\theta}_{k}=\underset{\theta}{\text{argmin}}~\frac{1}{N}\sum_{i=0}^{N-1}(f_{\theta}(\mathbf{s}_{i},\mathbf{a}_{i})-f_{\bar{\theta}_{k}}(\mathbf{s}_{i},\mathbf{a}_{i}))^{2}$ (2)
@@
To address these difficulties, we introduce a coupled approach. This approach jointly estimates both the expert distribution and the behavioral distribution; it encourages policy exploration during the early stages of training. We provide an intuitive illustration in Figure 2 and describe the detailed methodology in Section 3.3. In this coupled approach, we need to estimate the behavioral distribution during online training, which naturally raises the problem of inconsistent final rewards, as noted by [yang2024exploration]. Thus, we adopt their method for tracking the frequency of data occurrence, which we describe in Section 3.2.
@@ 3.2 Mitigating Inconsistent Reward Estimation
Inconsistencies can arise at various stages of RND model training [yang2024exploration]. During the initial stage, these inconsistencies stem from extreme values in one network, which can be mitigated by using multiple target networks (denoted as $K$ target networks). In the final stage, inconsistencies occur when the resulting reward distribution does not align with the actual state-action distribution. To address this, an unbiased estimator for the state-action occurrence count $n$ is necessary. We should track state-action occurrence frequencies in order to maintain consistency when the distributional RND model is trained online. In this section, we replace the original state $\mathbf{s}_{t}$ with the latent representation $\mathbf{z}_{t}$ for the input of the RND model. Following [yang2024exploration], we denote the random variable $c(\mathbf{z}_{t},\mathbf{a}_{t})$ as the output of a target network $f_{\bar{\theta}_{k}}$, where $k$ is sampled uniformly from the interval $[0,K)$. For a predictor $f$ estimating a distribution $\rho$ (which can be either the expert distribution $\rho_{E}$ or the behavioral distribution $\rho_{\pi}$), by minimizing the $L_{2}$-norm loss $\|f(\mathbf{z}_{t},\mathbf{a}_{t})-c(\mathbf{z}_{t},\mathbf{a}_{t})\|_{2}^{2}$, the optimal predictor $f^{*}(\mathbf{z}_{t},\mathbf{a}_{t})$ is given by:
@@
where $c_{i}(\mathbf{z}_{t},\mathbf{a}_{t})$ is representing the $c(\mathbf{z}_{t},\mathbf{a}_{t})$ for the $i$-th occurrence for state-action pair $(\mathbf{z}_{t},\mathbf{a}_{t})$ in distribution $\rho$. In order to track the occurrence count $n$, we adopt a lemma proposed by [yang2024exploration]:
@@
See Appendix E or prior work [yang2024exploration]. ∎
In this way, we are able to estimate the data distribution with higher consistency as the training proceeds. Following [yang2024exploration], we construct the following estimator for $\sqrt{1/n}$ as an additional bonus correction term:
@@ 3.3 Coupled Distributional Random Expert Distillation
We construct a reward model with two predictor networks that share the same random target ensemble on the latent space of a world model. The distributional random target ensemble consists of $K$ random networks $\{f_{\bar{\theta}_{k}}\}_{0:K}$ with fixed parameters. Regarding the predictors, one of them is the expert predictor $f_{\phi}$ while the other is the behavioral predictor $f_{\psi}$. A predictor $f$ is defined by $f:\mathcal{Z}\times\mathcal{A}\rightarrow\mathbb{R}^{p}$, while $p$ is the dimension of the low-dimensional embedding space for L2 norm distance computation. Following [yang2024exploration], we ask these two predictors to learn the random targets sampled. This is different to RED which learn $K$ predictors for $K$ targets. Given an expert buffer $\mathcal{B}_{E}$ and a behavioral buffer $\mathcal{B}_{\pi}$, we aim to optimize through the following objective:
@@
The scalar coefficient $\alpha$ in Eq. 8 balances the contributions of the first term (the $L_{2}$-norm) and the second term (an estimator for $\sqrt{1/n}$). Following [yang2024exploration], we let the first term dominate initially, switching to the second term as training progresses. This can be achieved with a fixed $\alpha$, rather than a dynamic coefficient. This modification enables consistent online estimation of the state-action distribution, directly supporting reward modeling for online imitation learning.
@@ 3.4 Integrating into World Models for Imitation Learning
World models learn the policy and underlying environment dynamics by encoding the observations into a latent space and learning the transition model in the latent space. Decoder-free world models such as TD-MPC series [hansen2022temporal, hansen2023td] has proved to be a powerful tool for complex reinforcement learning tasks. We leverage a decoder-free world model containing the following components:
@@
The first term contains consistency loss and temporal difference loss to ensure the prediction consistency of the dynamics model and the accuracy for value function estimation. the temporal difference target is computed by $q_{t}=R(\mathbf{z}_{t},\mathbf{a}_{t})+\gamma Q(\mathbf{z}^{\prime}_{t},\pi(\mathbf{z}^{\prime}_{t}))$ where $R(\mathbf{z}_{t},\mathbf{a}_{t})$ is the output of the reward model. We convert the regression TD objective into a classification problem for stabler value estimation, which is also used by the TD-MPC series and mentioned by [farebrother2024stop]. $\text{CE}(\hat{q}_{t},q_{t})$ is the cross entropy loss between target Q value and current predicted value. The second term is the reward loss, which is shown in Eq.6. Similar to the computation of reward loss, we also sum up the consistency and TD loss with factor $\lambda$ over a horizon $H$.
@@
Regarding the policy prior update, we adopt maximum entropy objective [haarnoja2018soft] to train a stochastic policy:
@@
Following TD-MPC series [hansen2022temporal, hansen2023td], we also leverage model predictive path integral (MPPI) [williams2015model] for planning. We optimize using the sampled action sequences $(\mathbf{a}_{t},\mathbf{a}_{t+1},...,\mathbf{a}_{t+H})$ in a derivative-free style, maximizing the estimated return for the latent trajectories that have been rolled out using our dynamics model. Mathematically, our objective can be describe as a return maximizing process [hansen2023td]:
@@ 4 Experiments
We conduct experiments across a diverse range of tasks, including locomotion, manipulation, and tasks with both visual and state-based observations. We evaluate our approach using the DMControl [tassa2018deepmind], Meta-World [yu2020meta] and ManiSkill2 [gu2023maniskill2] environments. As for the baselines, we compare our approach with IQ-MPC [li2024rewardfreeworldmodelsonline], which integrates a world model architecture, as well as with model-free approaches, specifically IQ-Learn+SAC [garg2021iq] (referred to as IQL+SAC in the plots), CFIL+SAC [freund2023coupled], HyPE [ren2024hybrid] (In Appendix D.3) and SAIL [SAIL] (In Appendix D.4). Additionally, we also incorporate behavioral cloning (BC) as a baseline method in our evaluation. For all experiments, we sample expert trajectories from a trained TD-MPC2 [hansen2023td]. Additionally, we conduct ablation studies on the number of expert trajectories and the choice of function $g$, as detailed in Appendix D.2. We further evaluate the robustness of our algorithm in noisy environment dynamics (Appendix D.5), examine the benefits of constructing the reward model in the latent space (Appendix D.8), and highlight its advantages over existing adversarial training methods (Appendix D.7). We also provide the quantitative results measuring the training stability in Appendix D.6. All of the experiments are conducted on a single RTX3090 graphic card.
@@ 4.1 Meta-World Experiments
We conduct experiments on 6 tasks in Meta-World environments. We use 100 expert trajectories for each task, ensuring that the expert data remains consistent across all algorithms for fair comparison within each task. IQ-MPC suffers from overly powerful discriminators in these tasks, even with gradient penalty applied, due to the adversarial training methodology. CFIL+SAC [freund2023coupled] encounters instability in the training process due to the challenges inherent in training flow models. We show stable and expert-level performance, outperforming these baselines in these tasks. We show the episode reward results in Figure 4 and success rate results in Table 1.
@@
[…]